OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (campbell_st):

can you find the slope of the parallel line

OpenStudy (anonymous):

y = mx + b m = slope of parallel line

OpenStudy (campbell_st):

here is the diagram 

OpenStudy (campbell_st):

so now you have the diagram do you think you can find the slope..?

OpenStudy (anonymous):

so i just divide 15/5 and get 3 right?

OpenStudy (campbell_st):

thats right so the equation of the line is your line will be y = 3x + b you have the point x = -2, y = 3 substitute then into the equation above for find b.. does that make sense

OpenStudy (anonymous):

yes so i would take y = 3x + b and plug in -2 and 3 into the equation and i get 3=-6+b???